# Coning off totally geodesic boundary components of a hyperbolic manifold

### v1

@@ Abstract.
Group-theoretically, we conclude that the fundamental group of the cone-off is hyperbolic and the $\pi_{1}$-image of the coned-off locally convex subset is a quasi-convex subgroup.
@@ 1. Introduction
The idea of coning off ends or boundary components of an aspherical manifold to obtain an aspherical space goes back at least to a comment of Gromov in [Gro93, 7.A.VI]. This was fleshed out in the non-compact real hyperbolic case by Mosher and Sageev [MS97]. Mosher and Sageev begin with a hyperbolic manifold with cusps. Under the condition that disjoint cusp cross sections with large (greater than $\pi$) injectivity radius can be chosen, they show that a negatively curved space can be obtained by cutting off the cusps and replacing them with cones. The method is differential geometric, utilizing the same kind of warped product construction as in the Gromov-Thurston $2\pi$ theorem [BH96]. Modifying the construction, a wide variety of compact non-positively curved spaces can be obtained by replacing the cusps with mapping cylinders of fibrations [FM10].
@@
We also give conditions of the same flavor under which a closed locally convex subset of the original manifold gives a locally convex (and hence $\pi_{1}$–injective) subset of the coned-off space. These results are used in work of the second author with Ruffoni to give examples of negatively curved three-dimensional pseudo-manifolds whose fundamental groups do not act properly on any CAT$(0)$ cube complex [MR].
As a consequence of our geometric construction, we obtain group-theoretic conclusions about the fundamental groups of the coned-off spaces and of the coned-off locally convex subsets. These are quantitative versions of results in “group-theoretic Dehn filling” [GM08, Osi07, GM21].
@@ 1.1. Statement of main theorems
[…]
@@
If $S\subset M$, we define $\widehat{S}\subset\widehat{M}$ to be the union of $S$ with the set of equivalence classes of points of the form $(s,t)$ where $s\in S$, $t\in[0,1]$.
@@
For $N$ a non-positively curved manifold without boundary, and $x\in N$, the buffer width $\operatorname{BW}_{N}(x)$ is equal to the injectivity radius at $x$, usually written $\operatorname{inj}(x)$. If $Y\subset N$ is a totally geodesic submanifold, then $\operatorname{BW}_{X}(Y)$ is the normal injectivity radius of $Y$.
@@
Here is the first main result. (Recall $\operatorname{inj}(X)=\inf_{x\in X}\operatorname{inj}(x)$.)
@@
Let $M$ be a hyperbolic manifold with totally geodesic boundary and suppose there are constants $b$ and $c$ so that
@@
Then there is a negatively curved metric $\hat{d}$ on $\widehat{M}$ and a locally isometric embedding of $M\smallsetminus N_{b}(\partial M)$ into $(\widehat{M},\hat{d})$ with image equal to $M\subset\widehat{M}$.
@@
Let $M_{0}$ be a compact hyperbolic $n$–manifold with totally geodesic boundary. There is a finite-sheeted cover $M\to M_{0}$ so that $\widehat{M}$ admits a negatively curved metric.
@@
Let $b=\frac{1}{2}\operatorname{BW}_{M_{0}}(\partial M_{0})$, and note that for any finite-sheeted cover $M\to M_{0}$, we have $\operatorname{BW}_{M}(\partial M)=\operatorname{BW}_{M_{0}}(\partial M_{0})$. There are finitely many conjugacy classes $c_{1},\ldots,c_{k}$ in $\pi_{1}M_{0}$ representing nontrivial loops of length at most $2\pi/\sinh(b)$ in $\partial M_{0}$. Since $\pi_{1}M_{0}$ is linear, it is residually finite [Mal40]. In particular there is a finite index normal subgroup of $\pi_{1}M_{0}$ missing the conjugacy classes $c_{1},\ldots,c_{k}$. The corresponding finite-sheeted cover $M\to M_{0}$ satisfies $\operatorname{inj}(\partial M)>\pi/\sinh(b)$. ∎
@@
Let $M$ be hyperbolic with totally geodesic boundary, and suppose $S\subset M$ is a closed locally convex set. For $t\in[0,\operatorname{BW}_{M}(\partial M))$, let $P_{t}=P_{t}(S)$, as in Definition 1.7. Suppose that $b,c>0$ satisfy the conditions (A1) and (A2) of Theorem A and additionally that there is a $b^{\prime}\in(b,\operatorname{BW}_{M}(\partial M))$ so that the following hold.
@@
    There is an isotopy $\Phi\colon\thinspace[0,b^{\prime}]\times\partial M\to\partial M$ so that $\Phi_{0}$ is the identity and $\Phi_{t}(P_{0})=P_{t}$ for all $t\in[0,b^{\prime}]$.
@@
The metric in Theorem B can be obtained from the metric in Theorem A via an isotopy.
@@
The hypotheses of Theorem B may be clarified by considering the situation of a properly embedded totally geodesic submanifold $S$ meeting the boundary of $M$. For appropriately chosen $b,b^{\prime}$, $S$ will always satisfy Condition 1. However, $S$ only satisfies hypothesis 2 when it meets the boundary orthogonally, in which case $P_{t}=P_{0}$ for small $t$. (If $S$ does not meet the boundary orthogonally, there may nonetheless be a “thickening” which satisfies hypothesis 2.) Condition 3 can often be ensured in a finite-sheeted cover, as the next Corollary shows.
@@
Let $M_{0}$ and $b,c>0$ satisfy the conditions of Theorem A, and let $S_{0}\subset M_{0}$ and $b^{\prime}>0$ satisfy all the conditions of Theorem B except possibly Condition 3. Suppose further that $S_{0}$ is connected and $\pi_{1}S_{0}$ is separable in $\pi_{1}M_{0}$. Then there is a finite-sheeted cover $M\to M_{0}$ so that $S$ lifts to $S\subset M$ and a negatively curved metric $\hat{d}$ on $\widehat{M}$ so that $\widehat{S}$ is locally convex in $\widehat{M}$.
@@
Scott’s characterization of subgroup separability [Sco78] means that any compact subset $K$ of the cover $\tilde{M}_{0}$ of $M_{0}$ corresponding to $\pi_{1}S_{0}$ embeds in a finite-sheeted cover $M$ of $M_{0}$. The locally convex set $S_{0}$ lifts to a set $\tilde{S}_{0}$ in $\tilde{M}_{0}$ with infinite buffer width. Taking $K$ to be a closed $c$–neighborhood of $\tilde{S}_{0}$, the corresponding finite cover $M$ contains a lifted copy $S$ of $S_{0}$, which moreover satisfies condition 3. ∎
We remark that the condition of separability is always true if $n\leq 3$ by results of Scott [Sco78] and Wise [Wis21].
@@
Suppose that $M$ is a compact hyperbolic $n$–manifold with totally geodesic boundary satisfying the conditions of Theorem A. Let $\Gamma=\pi_{1}M$, and let $K\lhd\Gamma$ be the subgroup normally generated by loops in $\partial M$. Then $\Gamma/K$ is torsion-free, hyperbolic, and has cohomological dimension $n$.
Suppose in addition that $H<\Gamma$ is the fundamental group of a locally convex subset $S\subset M$ satisfying the conditions of Theorem B. Then the image of $H$ in $\Gamma/K$ is quasi-convex.
@@
The first part of the Corollary is immediate from Theorem A. If each component of $\partial M$ intersects $S$ either in a connected set or in the empty set, the second part is immediate from Theorem B, since the image $\bar{H}$ of $H$ in $\Gamma/K$ can be identified with the fundamental group of $\widehat{S}$. Since $\widehat{M}$ is negatively curved and $\widehat{S}$ is a locally convex subset, the fundamental group of $\widehat{S}$ injects into that of $\widehat{M}$ (see [BH99, II.4.14]). If there are components of $\partial M$ whose intersection with $S$ is disconnected, then $\pi_{1}\widehat{S}$ is still a quasi-convex subgroup, but is not equal to $\bar{H}$; rather, it contains $\bar{H}$ as a free factor. But a free factor of a group is always quasi-isometrically embedded, so the quasi-convexity of $\bar{H}$ follows from the quasi-convexity of $\pi_{1}\widehat{S}$. ∎
The first part of Corollary 1.11 is related to the Relatively Hyperbolic Dehn Filling Theorem [GM08, Osi07] – the statement about cohomological dimension is related to work of Sun and Petrosyan [PS24]. The second part is connected to [GM21, Proposition 2.5]. All those statements though are for “sufficiently large” or “sufficiently $H$–wide” fillings. In principle the bounds for those results can be made explicit, but will not be as tight as the ones which we have provided.
@@ Acknowledgments
We would like to thank Daniel Groves for helpful conversations related to the correct statement of Theorem B. CK was supported by the NSF under grant DGE-1650441. JM was partially supported by the Simons Foundation, grant number 942496.
@@ 2.1. Warped products
We recall here the definition of a warped product of length spaces, which generalizes a warped product of Riemannian manifolds. For more information see [BBI01, §3.6.4].
@@
Let $(X,d_{X})$ and $(Y,d_{Y})$ be complete length spaces and let $f\colon\thinspace X\to[0,\infty)$ be a continuous function. For a lipschitz path $\gamma\colon\thinspace[a,b]\to X\times Y$, write $\gamma_{X}$ and $\gamma_{Y}$ for the projections to $X$ and $Y$ respectively. These are also lipschitz paths so their speeds $|\gamma_{X}^{\prime}|$, $|\gamma_{Y}^{\prime}|$ are defined almost everywhere [BBI01, Theorem 2.7.6]. Define a length
[…]
and use this to define a length pseudo-metric $d_{f}$. The warped product $X\times_{f}Y$ is the canonical metric space quotient of $(X\times Y,d_{f})$. (This is homeomorphic to $X\times Y$ if $f$ has no zeroes; otherwise some copies of $Y$ are smashed to points.)
@@
For example, the warping function $f\equiv 1$ yields the standard product metric. More interestingly, hyperbolic space $\mathbb{H}^{n+1}$ is isometric to a warped product in many ways, for example ×$\mathbb{R}~{}\times_{e^{t}}~{}\mathbb{E}^{n}$ or $[0,\infty)\times_{\sinh(t)}\mathbb{S}^{n}$. Most relevant for our purposes is the isometry
@@
For a compact hyperbolic manifold $M$ with totally geodesic boundary, and $r\leq\operatorname{BW}_{M}(\partial M)$, there is a Riemannian isometry
@@ 2.2. Local convexity
We first show the function $\mathrm{lcr}_{S}:S\to\mathbb{R}$ is lower semi-continuous. Fix $s\in S$ and $\varepsilon>0.$ Without loss of generality, we may assume that
$\varepsilon<\min\{\operatorname{BW}_{Z}(s),\mathrm{lcr}_{S}(s)\}.$
We must show that there exists $\delta>0$ such that $\mathrm{lcr}_{S}(s)-\varepsilon<\mathrm{lcr}_{S}(c)$ for all $c$ in $B_{\delta}(s)\cap S$. Set $\delta=\frac{\varepsilon}{2}.$ Let $c$ be a point in $B_{\delta}(s)\cap S$. Any ball $B_{r}(c)$ which is contained in $B_{\mathrm{lcr}_{S}(s)}(s)$ will have convex intersection with $S$; this can be seen by lifting to the universal cover of $Z$ (which we may do because $\varepsilon<\operatorname{BW}_{Z}(s)$). In particular, since
$\delta=\frac{\varepsilon}{2}<\mathrm{lcr}_{S}(s)-d_{Z}(s,c),$
the intersection $B_{\delta}(c)\cap B_{\mathrm{lcr}_{S}(s)}(s)$ is convex in $Z$. The intersection of two convex sets is convex, so for such $r$, $B_{\mathrm{lcr}_{S}(s)}(s)\cap S\cap B_{\delta}(c)=S\cap B_{\delta}(c)$ is convex.
Now suppose by way of contradiction that there exists a sequence of points $(s_{i})\subset S$ such that $\mathrm{lcr}_{S}(\{s_{i}\})=0$. Since $S$ is compact, we may suppose that $(s_{i})$ converges to some $s_{\infty}\in S$. Since $\mathrm{lcr}_{S}:S\to\mathbb{R}$ is lower semi-continuous,
$0=\lim_{i\to\infty}\mathrm{lcr}_{S}(s_{i})\geq\mathrm{lcr}_{S}(s_{\infty}),$
but $\mathrm{lcr}_{S}(s_{\infty})>0$ since $s_{\infty}$ is in the locally convex set $S$. ∎

### v2

@@ Abstract.
Group-theoretically, we conclude that the fundamental group of the cone-off is hyperbolic and the $\pi_{1}$-image of the coned-off locally convex subset is a quasi-convex subgroup. We also describe the Gromov boundary of the resulting hyperbolic group.
@@ 1. Introduction
The idea of coning off ends or boundary components of an aspherical manifold to obtain an aspherical space goes back at least to a comment of Gromov in [Gromov93, 7.A.VI]. This was fleshed out in the non-compact real hyperbolic case by Mosher and Sageev [MosherSageev]. Mosher and Sageev begin with a hyperbolic manifold with cusps. Under the condition that disjoint cusp cross sections with large (greater than $\pi$) injectivity radius can be chosen, they show that a negatively curved space can be obtained by cutting off the cusps and replacing them with cones. The method is differential geometric, utilizing the same kind of warped product construction as in the Gromov-Thurston $2\pi$ theorem [BleilerHodgson]. Modifying the construction, a wide variety of compact non-positively curved spaces can be obtained by replacing the cusps with mapping cylinders of fibrations [FM]. (In this paper non-positively curved means locally $\operatorname{CAT}(0)$, whereas negatively curved means locally $\operatorname{CAT}(k)$ for some $k<0$.)
@@
We also give conditions of the same flavor under which a closed locally convex subset of the original manifold gives a locally convex (and hence $\pi_{1}$–injective) subset of the coned-off space. These results are used in work of the second author with Ruffoni to give examples of negatively curved three-dimensional pseudo-manifolds whose fundamental groups do not act properly on any $\operatorname{CAT}(0)$ cube complex [MR].
As a consequence of our geometric construction, we obtain group-theoretic conclusions about the fundamental groups of the coned-off spaces and of the coned-off locally convex subsets. These are quantitative versions of results in “group-theoretic Dehn filling” [GM, Osin, GMQC].
@@ 1.1. Statement of main theorems
($\ddagger$}) $\widehat{M}=M\sqcup\left(\partial M\times[0,1]\right)/\sim,$
@@
If $S\subset M$, we define $\widehat{S}\subset\widehat{M}$ to be the union of $S$ with the set of equivalence classes of points of the form $(s,\zeta)$ where $s\in S\cap\partial M$, $\zeta\in[0,1]$.
@@
For $N$ a complete non-positively curved Riemannian manifold without boundary, and $x\in N$, the buffer width $\operatorname{BW}_{N}(x)$ is equal to the injectivity radius at $x$, usually written $\operatorname{inj}(x)$. If $Y\subset N$ is a totally geodesic submanifold, then $\operatorname{BW}_{X}(Y)$ is the normal injectivity radius of $Y$.
@@
Before stating the main result we clarify some notation. First, if $B$ is any metric space, $r>0$ and $A\subseteq B$, then $N_{r}(A)$ is the open rr–neighborhood of $A$. Second, if $X$ is a Riemannian manifold, $\operatorname{inj}(X)=\inf_{x\in X}\operatorname{inj}(x)$.
@@
Let $M$ be a compact hyperbolic manifold with totally geodesic boundary and suppose there are constants $b$ and $c$ so that
@@
[…]
We describe the boundary at infinity of the universal cover of $\widehat{M}$ in Appendix A
@@
Let $M_{0}$ be a compact hyperbolic manifold with totally geodesic boundary. There is a finite-sheeted cover $M\to M_{0}$ so that $\widehat{M}$ admits a negatively curved metric.
@@
For any finite-sheeted cover $M\to M_{0}$, we have $\operatorname{BW}_{M}(\partial M)=\operatorname{BW}_{M_{0}}(\partial M_{0})$. Indeed any geodesic segment orthogonal to the boundary in $M_{0}$ lifts to one in $M$; conversely any geodesic segment orthogonal to the boundary in $M$ projects to one in $M_{0}$.
There are finitely many conjugacy classes $c_{1},\ldots,c_{k}$ in $\pi_{1}M_{0}$ representing nontrivial loops of length at most $\pi/\sinh(\operatorname{BW}_{M}(\partial M))$ in $\partial M_{0}$. Since $\pi_{1}M_{0}$ is linear, it is residually finite [Malcev40]. In particular there is a finite index normal subgroup of $\pi_{1}M_{0}$ missing the conjugacy classes $c_{1},\ldots,c_{k}$. The corresponding finite-sheeted cover $M\to M_{0}$ satisfies $\operatorname{inj}(\partial M)>\pi/\sinh(\operatorname{BW}_{M}(\partial M))$. In particular Theorem A applies to $M$ by Remark 1.5. ∎
@@
Let $M$ be compact hyperbolic with totally geodesic boundary, and suppose $S\subset M$ is a closed locally convex set. For $t\in[0,\operatorname{BW}_{M}(\partial M))$, let $P_{t}=P_{t}(S)$, as in Definition 1.7. Suppose that $b,c>0$ satisfy the conditions (A1) and (A2) of Theorem A and additionally that there is a $b^{\prime}\in(b,\operatorname{BW}_{M}(\partial M))$ so that the following hold.
@@
    There is an isotopy $\Phi\colon\thinspace[0,b^{\prime}]\times\partial M\to\partial M$ so that $\Phi_{0}$ is the identity and $\Phi_{t}(P_{0})=P_{t}$ for all $t\in[0,b^{\prime}]$. [^2] Here we use the notation $\Phi_{t}(x)=\Phi(t,x)$.
@@
In the proof of Theorem B, the metric with respect to which $\widehat{S}$ is locally convex is obtained from the metric in Theorem A via an isotopy.
@@
The hypotheses of Theorem B may be clarified by considering the situation of an embedded totally geodesic submanifold $S$ so that $\partial S\subset\partial M$. For appropriately chosen $b,b^{\prime}$, $S$ will always satisfy Condition 1. However, $S$ only satisfies hypothesis 2 when it meets the boundary orthogonally (or not at all), in which case $P_{t}=P_{0}$ for small $t$. (If $S$ meets the boundary, but not orthogonally, there may nonetheless be a “thickening” which satisfies hypothesis 2.) Condition 3 can often be ensured in a finite-sheeted cover, as the next Corollary shows.
@@
Let $M_{0}$ and $b,c>0$ satisfy the conditions of Theorem A, and let $S_{0}\subset M_{0}$ and $b^{\prime}>0$ satisfy all the conditions of Theorem B except possibly Condition 3. Suppose further that $S_{0}$ is connected and $\pi_{1}S_{0}$ is separable in $\pi_{1}M_{0}$. Then there are a finite-sheeted cover $M\to M_{0}$ so that $S_{0}$ lifts (homeomorphically) to $S\subset M$ and a negatively curved metric $\hat{d}$ on $\widehat{M}$ so that $\widehat{S}$ is locally convex in $\widehat{M}$.
@@
Scott’s characterization of subgroup separability [Scott78] means that any compact subset $K$ of the cover $\widetilde{M}_{0}$ of $M_{0}$ corresponding to $\pi_{1}S_{0}$ embeds in a finite-sheeted cover $M$ of $M_{0}$. The locally convex set $S_{0}$ lifts to a set $S=\widetilde{S}_{0}$ in $\widetilde{M}_{0}$ with infinite buffer width. Taking $K$ to be a closed cc–neighborhood of $\widetilde{S}_{0}$, the corresponding finite cover $M$ contains a lifted copy $S$ of $S_{0}$, which moreover satisfies condition 3. ∎
We remark that the condition of separability is always true if $\dim(M)\leq 3$ by results of Scott [Scott78] and Wise [Wise21].
@@
Suppose that $M$ is a compact hyperbolic manifold with totally geodesic boundary satisfying the conditions of Theorem A. Let $\Gamma=\pi_{1}M$, and let $K\lhd\Gamma$ be the subgroup normally generated by loops in $\partial M$. Then $\Gamma/K$ is torsion-free, hyperbolic, and has cohomological dimension equal to $\dim(M)$.
Suppose in addition that $H<\Gamma$ is the fundamental group of a connected locally convex subset $S\subset M$ satisfying the conditions of Theorem B. Then the image of $H$ in $\Gamma/K$ is quasi-convex.
@@
The first part of the Corollary is immediate from Theorem A. If each component of $\partial M$ intersects $S$ either in a connected set or in the empty set, the second part is immediate from Theorem B, since the image $\bar{H}$ of $H$ in $\Gamma/K$ can be identified with the fundamental group of $\widehat{S}$. Since $\widehat{M}$ is negatively curved and $\widehat{S}$ is a locally convex subset, the fundamental group of $\widehat{S}$ injects into that of $\widehat{M}$ (see [BH, II.4.14]). If there are components of $\partial M$ whose intersection with $S$ is disconnected, then $\pi_{1}\widehat{S}$ is still a quasi-convex subgroup, but is not equal to $\bar{H}$; rather, it contains $\bar{H}$ as a free factor by van Kampen’s theorem. But a free factor of a group is always quasi-isometrically embedded, so the quasi-convexity of $\bar{H}$ follows from the quasi-convexity of $\pi_{1}\widehat{S}$. ∎
The first part of Corollary 1.11 is related to the Relatively Hyperbolic Dehn Filling Theorem [GM, Osin] – the statement about cohomological dimension is related to work of Sun and Petrosyan [SP]. The second part is connected to [GMQC, Proposition 2.5]. All those statements though are for “sufficiently large” or “sufficiently HH–wide” fillings. In principle the bounds for those results can be made explicit, but will not be as tight as the ones which we have provided.
@@ Acknowledgments
We would like to thank Daniel Groves for helpful conversations related to the correct statement of Theorem B. Thanks also to Lorenzo Ruffoni for helpful conversations and in particular asking the question which inspired this project. Thank you to Sam Taylor for a question that helped clarify the language of Remark 1.9. Finally we thank the referee for thoughtful comments and suggestions. CK was supported by the NSF under grant DGE-1650441. JM was partially supported by the Simons Foundation, grant number 942496.
@@ 2.1. Warped products
We recall here the definition of a warped product of length spaces, which generalizes a warped product of Riemannian manifolds. For more information see [BBI, §3.6.4].
@@
Let $(X,d_{X})$ and $(Y,d_{Y})$ be length spaces and let $f\colon\thinspace X\to[0,\infty)$ be a continuous function. For a Lipschitz path $\gamma\colon\thinspace[a,b]\to X\times Y$, write $\gamma_{X}$ and $\gamma_{Y}$ for the projections to $X$ and $Y$ respectively. These are also Lipschitz paths so their speeds $|\gamma_{X}^{\prime}|$, $|\gamma_{Y}^{\prime}|$ are defined almost everywhere [BBI, Theorem 2.7.6]. Define a length
$\ell_{f}(\gamma):=\int_{a}^{b}\sqrt{|\gamma^{\prime}_{X}(t)|^{2}+f^{2}(\gamma_{X}(t))|\gamma^{\prime}_{Y}(t)|^{2}}dt,$
and use this to define a length pseudo-metric $d_{f}$. The warped product $X\times_{f}Y$ is the canonical metric space quotient of $(X\times Y,d_{f})$. (This is homeomorphic to $X\times Y$ if $f$ has no zeroes; otherwise some copies of $Y$ are smashed to points.) The space $X$ is referred to as the base; the space $Y$ as the fiber.
@@
For example, the warping function $f\equiv 1$ yields the standard product metric. More interestingly, hyperbolic space $\mathbb{H}^{n+1}$ is isometric to a warped product in many ways, for example $\mathbb{R}~\times_{e^{t}}~\mathbb{E}^{n}$ or $0,\infty)\times_{\sinh(t)}\mathbb{S}^{n}$. Most relevant for our purposes is the isometry
@@
For a compact hyperbolic manifold $M$ with totally geodesic boundary, and $r\leq\operatorname{BW}_{M}(\partial M)$, there is an isometry
@@
We will largely be interested in examples in which the base is an interval and the warping function is increasing, convex, and has a unique zero. We record the following observations about geodesics in this setting.
Lemma 2.3.
Let $I\subset\mathbb{R}$ be an interval with $\inf(I)=t_{0}$, and let $f:I\to[0,\infty)$ be continuous and satisfy $f^{-1}(0)=\{t_{0}\}$. Consider $C=I\times_{f}F$ for $F$ a complete length space. Let $\ast$ be the cone point, ie the point of $C$ corresponding to the equivalence class $\{t_{0}\}\times F$. Let
$\pi_{I}\colon\thinspace C\to I,\quad\pi_{F}\colon\thinspace C\smallsetminus\{\ast\}\to F$
be the projections, and let $\sigma:[a,a^{\prime}]\to C$ be geodesic.
  1. (1)
    If $\sigma^{-1}(\ast)$ is nonempty, then $\sigma^{-1}(\ast)=\{s_{0}\}$ for some $s_{0}\in I$ and $\pi_{F}\circ\sigma|_{I\smallsetminus\{s_{0}\}}$ is locally constant.
If moreover $f$ is strictly increasing on $I$, then
  1. (2)
    the composition $\pi_{I}\circ\sigma$ is convex; and
  2. (3)
    if $\sigma$ misses the cone point, then $\pi_{F}\circ\sigma$ is a reparametrized geodesic.
If moreover $f$ is convex, then
  1. (4)
    the geodesic $\sigma$ passes through the cone point if and only if
    $d_{F}(\pi_{F}(\sigma(a)),\pi_{F}(\sigma(a^{\prime})))\geq\pi.$
@@ 2.2. Local convexity
Take a cover of $S$ by balls around all $x\in S$ of radius $\frac{r_{x}}{2}$ for where
$r_{x}=\min\{\mathrm{lcr}_{S}(x),\operatorname{BW}_{Z}(x)\}.$
Pass to a finite subcover and let $\varepsilon$ be the minimum $r_{x}$ appearing in the subcover. We aim to show that $\mathrm{lcr}_{S}(S)\geq\frac{\varepsilon}{2}$. Fix some $B_{\frac{r_{x}}{2}}(x)$ in the finite subcover, and for readability write $\frac{r_{x}}{2}=r_{0}$. Let $c$ be a point in $B_{r_{0}}(x)$. We shall show $S\cap B_{r_{0}}(c)$ is convex in $Z$. Observe:
$S\cap B_{r_{0}}(c)$ $=S\cap B_{r_{0}}(c)\cap B_{r_{x}}(x)$
$=\left(B_{r_{0}}(c)\cap B_{r_{x}}(x)\right)\cap\left(S\cap B_{r_{x}}(x)\right).$
We claim that this is an intersection of two convex sets, hence convex. Indeed, $S\cap B_{\mathrm{lcr}_{S}(s)}(s)$ is convex by definition. As $B_{r_{x}}(x)$ contains $B_{r_{0}}(c)$ and $r_{x}<\operatorname{BW}_{Z}(x)$, we can lift $B_{r_{x}}(x)$ to the universal cover of $Z$, which is $\operatorname{CAT}(0)$, so balls are convex, showing that $B_{r_{0}}(c)\cap B_{r_{x}}(x)$ is convex. Therefore,
$\mathrm{lcr}_{S}(c)\geq r_{0}\geq\frac{\varepsilon}{2}.$